OpenStudy (anonymous):

What is the Volume of a right cylinder with radius 2ft and heights 5ft? Leave answer in terms of pi please.

Directrix (directrix):

What is the formula?

OpenStudy (anonymous):

4/3 times the cube of the radius

Directrix (directrix):

If that's the formula, then how does the height of 5 feet fit into this?

Directrix (directrix):

4/3 times the cube of the radius --> I think this is the volume of a sphere formula.

OpenStudy (anonymous):

The volume of a cylinder equals the area of a base times the height of the cylinder

Directrix (directrix):

Okay, then.

Directrix (directrix):

The base of a right circular cylinder is a circle. That would be pi* r^2 for that and then multiply by h to get the volume. (see file attached above) V = pi * (r)^2 (h) @DustinMHayes Fill in for r and h and crank it out. I'll do the same and we will compare answers.

OpenStudy (anonymous):

10pi ft to the third power

Directrix (directrix):

I need to see your steps. I did not get 10 pi cubic feet. V = pi * (2)^2 (5) V = pi * 2 * 2 * 5 V = ?

Directrix (directrix):

@DustinMHayes And, the answer is .....

OpenStudy (anonymous):

20pi ft to thw third power

Directrix (directrix):

20 pi cubic feet is what I got, too.

OpenStudy (anonymous):

thank you for your help

Directrix (directrix):

Glad to assist. It was fun.